# Persistence of AR($1$) sequences with Rademacher innovations and linear mod $1$ transforms

### v2

@@ 5. Properties of the mapping $a\mapsto\lambda_{a}$
We first state the properties of the trajectory of 00 under $T_{a}$ as a function of parameter $a$.
@@
We argue by induction. In the basis case $k=2$, we have $\{a\in[\frac{1}{2},\frac{2}{3}]:\varkappa_{a}\geq k\}=[\frac{1}{2},\frac{2}{3}]$, where both functions $T_{a}^{1}(0)=\frac{1}{a}$ and $T_{a}^{2}(0)=\frac{1-a}{a^{2}}$ are strictly decreasing and continuous. Therefore, since the range of the second one is $[\frac{3}{4},2]$ and the function $\frac{2a-1}{1-a}$ increases on $[\frac{1}{2},\frac{2}{3}]$ from 00 to $1$, we have
@@
and the sets under the union are disjoint, it follows from the assumption of induction for Part b) that $J=[a^{\prime},a^{\prime\prime}]$ for some distinct $a^{\prime}<a^{\prime\prime}$ that satisfy the following restrictions. If $\frac{1}{2}\in J$, then $a^{\prime}=\frac{1}{2}$ and $T_{a^{\prime\prime}}^{i}(0)=1$ for some integer $12\leq i\leq n-$ 1. If $\frac{2}{3}\in J$, then $a^{\prime\prime}=\frac{2}{3}$ and $T_{a^{\prime}}^{i}(0)=\frac{2a^{\prime}-1}{1-a^{\prime}}$ for some integer $12\leq i\leq n-$ 1. Otherwise, $T_{a^{\prime\prime}}^{i}(0)=1$ and $T_{a^{\prime}}^{j}(0)=\frac{2a^{\prime}-1}{1-a^{\prime}}$ for some integers $12\leq i,j\leq n-$ 1. Notice that in all these cases, we have
@@
Indeed, if $a^{\prime\prime}=\frac{2}{3}$, then the dichotomy above is trivial because in this case $\frac{2a^{\prime\prime}-1}{1-a^{\prime\prime}}=1$, otherwise the orbit of zero under $T_{a^{\prime\prime}}$ is purely periodic and thus 00 is its only point in $[0,1)$.
@@
Furthermore, let $J^{\prime}=J$ if $T^{n}_{a^{\prime\prime}}(0)\geq 1$, otherwise let $J^{\prime}$ be either of the intervals $[a^{\prime},b^{\prime\prime}]$ and $[b^{\prime},a^{\prime\prime}]$. Consider the mapping $(a,x)\mapsto T_{a}(x)$ defined on its domain $D$ given by
@@
This mapping is continuous on $D\setminus\{(\frac{2}{3},1)\}$. We have $(a,T_{a}^{n}(0))\in D\setminus\{(\frac{2}{3},1)\}$ for every $a\in J^{\prime}$ by the choice of $J^{\prime}$ and the fact that the orbit of 00 under $T_{2/3}$ does not hit $1$, as shown in Section 2.2. Therefore, $T_{a}^{n+1}(0)$ is continuous on $J^{\prime}$, as claimed, by $T_{a}^{n+1}(0)=T_{a}(T_{a}^{n}(0))$. Clearly, $\delta_{n}$ is constant on $J^{\prime}$, as claimed.
@@
Indeed, in the first inequality we used that $T^{n}_{a}(0)<T^{n}_{b}(0)$ by the assumption of induction and the facts that (i) $T_{a}(x)$ increases in $x$ on each of the intervals $[0,\frac{2a-1}{1-a})$ and $[1,\frac{1}{1-a}]$, and (ii) one of these intervals contains both quantities $T^{n}_{a}(0)$ and $T^{n}_{b}(0)$ by the choice of $J^{\prime}$. In the second inequality we used that $T_{a}(x)<T_{b}(x)$ for every fixed $x$ in the domain of $T_{b}$, which contains $T^{n}_{b}(0)$ (by the choice of $b$), and is included in the domain of $T_{a}$. ∎
We now prove the properties of $\lambda_{a}$ stated.
@@
Note in passing that it is easy to give a self-contained proof of the monotonicity using the facts that $\lambda_{a}$ is constant on every interval in $[\frac{1}{2},\frac{2}{3}]\setminus S$ and is constant on no open interval that meets $S$. We will prove these facts later on using equation (11).
@@
Recall that $S=\{a\in[1/2,2/3]:\varkappa_{a}=\infty\}$ and $\hat{T}_{a}(x)=\frac{1}{a}x+\frac{1}{2}\pmod{1}$ for $0\leq x\leq 1$. By Parry [19, Theorem 6] and Halfin [15, Theorem 4.4], the mapping $\hat{T}_{a}$ has an absolutely continuous invariant probability measure $\hat{\mu}_{a}$ on $[0,1]$ with the density proportional to
$\hat{h}_{a}(x)=\sum_{k=0}^{\infty}a^{k+1}\mathbbm{1}\{\hat{T}_{a}^{k}(0)\leq x\}-\sum_{k=0}^{\infty}a^{k+1}\mathbbm{1}\{\hat{T}_{a}^{k}(1)\leq x\}.$ (58)
Note in passing that there are no other invariant probability densities for $\hat{T}_{a}$.
It follows from (8) that $\varkappa_{a}=\inf\{k\geq 0:\hat{T}_{a}^{k}(0)\in\hat{I}_{a}\}$, where $\hat{I}_{a}=(\frac{a(2a-1)}{2(1-a)},\frac{a}{2})$. According to Corollary 1 in Faller and Pfister [13], the orbit of zero $\{\hat{T}_{a}^{k}(0):k\geq 0\}$ is $\hat{\mu}_{a}$-normal for almost all values of $a$. This means that for every continuous function $f$ on $[0,1]$, we have
$\lim_{n\to\infty}\frac{1}{n}\sum_{i=0}^{n-1}f(\hat{T}_{a}^{k}(0))=\int_{0}^{1}f(x)\hat{\mu}_{a}(dx).$
On the other hand, for $a\in(\frac{1}{2},\frac{2}{3})$ we have $\hat{\mu}_{a}(\hat{I}_{a})>0$ by Theorem 3 in Hofbauer [17], which asserts that the support of $\hat{\mu}_{a}$ is the whole of the interval $[0,1]$. Hence $\varkappa_{a}$ is finite for almost all $a\in(\frac{1}{2},\frac{2}{3})$, and thus $S$ has Lebesgue measure zero.
The set $S$ is closed because its complement in $[\frac{1}{2},\frac{2}{3}]$ is open by Lemma 14.b. The topological support of $d\lambda_{a}$ is a subset of $S$ since $\lambda_{a}$ is constant on every interval contained in $[\frac{1}{2},\frac{2}{3}]\setminus S$, as we proved above in Item 2. Then $d\lambda_{a}$ is singular with respect to the Lebesgue measure. It remains to show that $\lambda_{a}$ is constant on no open interval that meets $S$.
@@
because $\delta_{k}^{\prime}=0$ for all $k>k_{0}$. At least one term in the sum is strictly positive because $1\leq T_{a}^{k_{0}+1}(0)<\frac{1}{1-a}$, hence the trajectory of 00 eventually returns to $[0,1)$. This is a contradiction.
Now assume that $a^{\prime}\in S$ and $\{T_{a^{\prime}}^{k}(0):k\geq 1\}$ does not hit the point $\frac{2a^{\prime}-1}{1-a^{\prime}}$. Then $a^{\prime}>\frac{1}{2}$, and since the set $S$ has measure 00, there exists an increasing sequence $\{a_{n}\}_{n}\subset(1/2,a^{\prime})\setminus S$ that converges to $a^{\prime}$ as $n\to\infty$. Denote by $a_{n}^{\prime}$ the right endpoint of the maximal open interval in $[\frac{1}{2},\frac{2}{3}]\setminus S$ that contains $a_{n}$. The sequence $a_{n}^{\prime}$ increases and converges to $a^{\prime}$. We have $a_{n}^{\prime}<a^{\prime}$ for every $n$ because otherwise $T_{a^{\prime}}^{k}(0)=\frac{2a^{\prime}-1}{1-a^{\prime}}$ for some $k\geq 1$ by Lemma 14.b. Then $\lambda_{a_{n}^{\prime}}<\lambda_{a^{\prime}}$, as shown above, hence $\lambda_{a}<\lambda_{a^{\prime}}$ for all $a<a^{\prime}$. Thus, we showed that $\lambda_{a}$ is constant on no open neighbourhood of $a^{\prime}$.
4. The equality $\lambda_{1/2}(p)=p$ for all $p\in(0,1)$ is already established in Remark 2.i.
6. Convergence of the conditional distributions
In this section we will prove convergence of conditional distributions towards the quasi-stationary distribution $\nu_{a}$ and study the properties of this limiting distribution.
1. Convergence of conditional distributions.
The functional space $U$, which has been used in our analysis of the probabilities $\mathbb{P}_{x}(\tau>n)$, is quite narrow. This space does not contain indicator functions of all subintervals of $[0,1/(1-a)]$ and therefore, we cannot use $U$ to study the distribution of $X_{n}$ conditioned on $\{\tau>n\}$. For this reason, we shall now regard $P$ as an operator acting on the larger space $BV$ of functions of bounded variation on $[0,1/(1-a)]$ equipped with the standard norm (28).
@@ 6. Convergence of the conditional distributions
Then $P=P_{1}+P_{2}$ and $P_{1}f\in U$ for every function $f$ of bounded variation.
@@
$P^{n-j-1}(P_{1}P_{2}^{j}f)(x)\sim c(M(P_{1}P_{2}^{j}f),v^{*})\lambda_{a}^{n-j-1}V(x)$ (60)
for every fixed $j\geq 0$.
@@
$P_{2}f_{z}(x)=p\big{(}\mathbbm{1}\{ax+1>z\})-\mathbbm{1}\{1>z\}\big{)}+q{\mathbbm{1}}\{x\geq 1/a\}\big{(}\mathbbm{1}\{ax-1>z\})-\mathbbm{1}\{0>z\}\big{)}.$
@@
$P_{1}P_{2}^{j}f_{z}(x)=pP_{2}^{j}f_{z}(1)\quad\text{for every }x\in[0,1/(1-a)],$
@@
$\sum_{j=0}^{\infty}\lambda_{a}^{-j-1}P_{1}P_{2}^{j}f_{z}(x)=\sum_{j=0}^{\varkappa_{a}(z)}\delta_{j}(z)(p/\lambda_{a})^{j+1}(q/p)^{L_{j}(z)}=:\overline{F}_{a}(z).$
This sum converges for infinite $\varkappa_{a}(z)$ by $\lambda_{a}>p$ if $p\geq 1/2$ and by Proposition 8 if $p<1/2$.
Combining the above equalities with (59) and using asymptotics (60), where the rate of convergence is estimated by (34), we get
$P^{n}f_{z}(x)\sim c\overline{F}_{a}(z)\lambda_{a}^{n}V(x)$ (62)
as $n\to\infty$ for every $x,z\in[0,1/(1-a)]$ together with the bound
$|\lambda_{a}^{-n}P^{n}f_{z}(x)-c\overline{F}_{a}(z)V(x)|$ $\leq\lambda_{a}^{-n}P_{2}^{n}f_{z}(x)+C_{1}\gamma^{n}\sum_{j=0}^{n-1}\lambda_{a}^{-j-1}\|P_{1}P_{2}^{j}f_{z}\|$
$\qquad+\sum_{j=n}^{\infty}\lambda_{a}^{-j-1}P_{1}P_{2}^{j}f_{z}(x)$
$\leq C_{1}\gamma^{n}\overline{F}_{a}(z)+\sum_{j=n}^{\varkappa_{a}(z)}(p/\lambda_{a})^{j}(q/p)^{L_{j}(z)}.$
By Proposition 8 and the inequality $\overline{F}_{a}(z)\leq\overline{F}_{a}(0)$, this implies that
@@
for all $n\geq 0$ with some $C_{2}>0$ and $\gamma_{1}\in(0,1)$ given by $\gamma_{1}=\max(\gamma,(p/\lambda_{a})(q/p)^{C})$, where both constants $\gamma_{1}$ and $C_{2}$ depend only on $a$ and $p$.
@@
$\lim_{n\to\infty}\mathbb{P}_{x}(X_{n}>z|\tau>n)=\overline{F}_{a}(z)$
for every $x,z\in[0,1/(1-a)]$. This limit does not depend on the starting point $x$. It also follows that the function $\overline{F}_{a}$ is non-increasing. Therefore, the conditional distributions converge weakly to the measure $\nu_{a}$ on $[0,\frac{1}{1-a}]$ such that
@@
whenever $\overline{F}_{a}$ is continuous at $z$. This measure is a probability since the conditional distributions are supported on $[0,1/(1-a)]$ and therefore tight.
2. The support and non-atomicity of the measure $\nu_{a}$.

### v3

@@ 5. Properties of the mapping $a\mapsto\lambda_{a}$
In this section we prove the properties of $\lambda_{a}$ stated in Theorem 1. We first study the properties of the trajectory of 0 under $T_{a}$ as a function of parameter $a$.
@@
We argue by induction. In the basis case $k=2$, we have $\{a\in[\frac{1}{2},\frac{2}{3}]:\varkappa_{a}\geq k\}=[\frac{1}{2},\frac{2}{3}]$, where both functions $T_{a}^{1}(0)=\frac{1}{a}$ and $T_{a}^{2}(0)=\frac{1-a}{a^{2}}$ are strictly decreasing and continuous. Therefore, since the range of the second one is $\frac{3}{4},2]$ and the function $\frac{2a-1}{1-a}$ increases on $\frac{1}{2},\frac{2}{3}]$ from 0 to $1$, we have
@@
and the sets under the union are disjoint, it follows from the assumption of induction for Part b) that $J=[a^{\prime},a^{\prime\prime}]$ for some distinct $a^{\prime}<a^{\prime\prime}$ that satisfy the following restrictions. If $\frac{1}{2}\in J$, then $a^{\prime}=\frac{1}{2}$ and $T_{a^{\prime\prime}}^{i}(0)=1$ for some integer $2\leq i\leq n-1$. If $\frac{2}{3}\in J$, then $a^{\prime\prime}=\frac{2}{3}$ and $T_{a^{\prime}}^{i}(0)=\frac{2a^{\prime}-1}{1-a^{\prime}}$ for some integer $2\leq i\leq n-1$. Otherwise, $T_{a^{\prime\prime}}^{i}(0)=1$ and $T_{a^{\prime}}^{j}(0)=\frac{2a^{\prime}-1}{1-a^{\prime}}$ for some integers $2\leq i,j\leq n-1$. Notice that in all these cases, we have
@@
Indeed, if $a^{\prime\prime}=\frac{2}{3}$, then the dichotomy above is trivial because in this case $\frac{2a^{\prime\prime}-1}{1-a^{\prime\prime}}=1$, otherwise the orbit of zero under $T_{a^{\prime\prime}}$ is purely periodic and thus 0 is its only point in $[0,1)$.
@@
Furthermore, let $J^{\prime}=J$ if $T^{n}_{a^{\prime\prime}}(0)\geq 1$, otherwise let $J^{\prime}$ be either of the intervals $a^{\prime},b^{\prime\prime}]$ and $b^{\prime},a^{\prime\prime}]$. Consider the mapping $(a,x)\mapsto T_{a}(x)$ defined on its domain $D$ given by
@@
This mapping is continuous on $D\setminus\{(\frac{2}{3},1)\}$. We have $(a,T_{a}^{n}(0))\in D\setminus\{(\frac{2}{3},1)\}$ for every $a\in J^{\prime}$ by the choice of $J^{\prime}$ and the fact that the orbit of 0 under $T_{2/3}$ does not hit $1$, as shown in Section 2.3. Therefore, $T_{a}^{n+1}(0)$ is continuous on $J^{\prime}$, as claimed, by $T_{a}^{n+1}(0)=T_{a}(T_{a}^{n}(0))$. Clearly, $\delta_{n}$ is constant on $J^{\prime}$, as claimed.
@@
Indeed, in the first inequality we used that $T^{n}_{a}(0)<T^{n}_{b}(0)$ by the assumption of induction and the facts that (i) $T_{a}(x)$ increases in $x$ on each of the intervals $0,\frac{2a-1}{1-a})$ and $1,\frac{1}{1-a}]$, and (ii) one of these intervals contains both quantities $T^{n}_{a}(0)$ and $T^{n}_{b}(0)$ by the choice of $J^{\prime}$. In the second inequality we used that $T_{a}(x)<T_{b}(x)$ for every fixed $x$ in the domain of $T_{b}$, which contains $T^{n}_{b}(0)$ (by the choice of $b$), and is included in the domain of $T_{a}$. ∎
We now prove the properties of $\lambda_{a}$ stated in Theorem 1.
@@
Note in passing that it is easy to give a self-contained proof of the monotonicity using the facts that $\lambda_{a}$ is constant on every interval in $\frac{1}{2},\frac{2}{3}]\setminus S$ and is constant on no open interval that meets $S$. We will prove these facts later on using equation (15).
@@
Recall that $S=\{a\in[1/2,2/3]:\varkappa_{a}=\infty\}$ and $\widehat{T}_{a}(x)=\frac{1}{a}x+\frac{1}{2}\pmod{1}$ for $0\leq x\leq 1$. By Parry [21, Theorem 6] and Halfin [16, Theorem 4.4], the mapping $\widehat{T}_{a}$ has an absolutely continuous invariant probability measure $\widehat{\mu}_{a}$ on $[0,1]$ with the density proportional to
$\widehat{h}_{a}(x)=\sum_{k=0}^{\infty}a^{k+1}\mathbbm{1}\{\widehat{T}_{a}^{k}(0)\leq x\}-\sum_{k=0}^{\infty}a^{k+1}\mathbbm{1}\{\widehat{T}_{a}^{k}(1)\leq x\}.$ (68)
Note in passing that there are no other invariant probability densities for $\widehat{T}_{a}$.
It follows from (10) that $\varkappa_{a}=\inf\{k\geq 0:\widehat{T}_{a}^{k}(0)\in\widehat{I}_{a}\}$, where $\widehat{I}_{a}=(\frac{a(2a-1)}{2(1-a)},\frac{a}{2})$. According to Corollary 1 in Faller and Pfister [14], the orbit of zero $\{\widehat{T}_{a}^{k}(0):k\geq 0\}$ is $\widehat{\mu}_{a}$-normal for almost all values of $a$. This means that for every continuous function $f$ on $[0,1]$, we have
$\lim_{n\to\infty}\frac{1}{n}\sum_{i=0}^{n-1}f(\widehat{T}_{a}^{k}(0))=\int_{0}^{1}f(x)\widehat{\mu}_{a}(dx).$
On the other hand, for $a\in(\frac{1}{2},\frac{2}{3})$ we have $\widehat{\mu}_{a}(\widehat{I}_{a})>0$ by Theorem 3 in Hofbauer [18], which asserts that the support of $\widehat{\mu}_{a}$ is the whole of the interval $[0,1]$. Hence $\varkappa_{a}$ is finite for almost all $a\in(\frac{1}{2},\frac{2}{3})$, and thus $S$ has Lebesgue measure zero.
The set $S$ is closed because its complement in $\frac{1}{2},\frac{2}{3}]$ is open by Lemma 16.b. The topological support of $d\lambda_{a}$ is a subset of $S$ since $\lambda_{a}$ is constant on every interval contained in $\frac{1}{2},\frac{2}{3}]\setminus S$, as we proved above in Item 2. Then $d\lambda_{a}$ is singular with respect to the Lebesgue measure. It remains to show that $\lambda_{a}$ is constant on no open interval that meets $S$.
@@
because $\delta_{k}^{\prime}=0$ for all $k>k_{0}$. At least one term in the sum is strictly positive because $1\leq T_{a}^{k_{0}+1}(0)<\frac{1}{1-a}$, hence the trajectory of 0 eventually returns to $[0,1)$. This is a contradiction.
Now assume that $a^{\prime}\in S$ and $\{T_{a^{\prime}}^{k}(0):k\geq 1\}$ does not hit the point $\frac{2a^{\prime}-1}{1-a^{\prime}}$. Then $a^{\prime}>\frac{1}{2}$, and since the set $S$ has measure 0, there exists an increasing sequence $\{a_{n}\}_{n}\subset(1/2,a^{\prime})\setminus S$ that converges to $a^{\prime}$ as $n\to\infty$. Denote by $a_{n}^{\prime}$ the right endpoint of the maximal open interval in $\frac{1}{2},\frac{2}{3}]\setminus S$ that contains $a_{n}$. The sequence $a_{n}^{\prime}$ increases and converges to $a^{\prime}$. We have $a_{n}^{\prime}<a^{\prime}$ for every $n$ because otherwise $T_{a^{\prime}}^{k}(0)=\frac{2a^{\prime}-1}{1-a^{\prime}}$ for some $k\geq 1$ by Lemma 16.b. Then $\lambda_{a_{n}^{\prime}}<\lambda_{a^{\prime}}$, as shown above, hence $\lambda_{a}<\lambda_{a^{\prime}}$ for all $a<a^{\prime}$. Thus, we showed that $\lambda_{a}$ is constant on no open neighbourhood of $a^{\prime}$.
4. The equality $\lambda_{1/2}(p)=p$ for all $p\in(0,1)$ is already established in (19).
6. Convergence to the quasi-stationary distribution $\nu_{a}$ and its properties
In this section we prove convergence of the conditional distributions stated in (17). Then we prove the properties of the limiting quasi-stationary distribution $\nu_{a}$ stated in Theorem 1.
6.1. Convergence of the conditional distributions
The functional space $U$, which we used in our analysis of the probabilities $\mathbb{P}_{x}(\tau>n)$, is quite narrow. This space does not contain indicator functions of all subintervals of [0,1/(1a)][0,1/(1-a)] and therefore, we cannot use $U$ to study the distribution of $X_{n}$ conditioned on $\{\tau>n\}$. For this reason, we shall now regard $P$ as an operator acting on the larger space $BV$ of functions of bounded variation on [0,1/(1a)][0,1/(1-a)] equipped with the standard norm (37).
@@ 6.1. Convergence of the conditional distributions
Then $P=P_{1}+P_{2}$ and $P_{1}f\in U$ for every function $f$ of bounded variation on [0,1/(1a)][0,1/(1-a)].
@@
$P^{n-j-1}(P_{1}P_{2}^{j}f)(x)\sim c(M(P_{1}P_{2}^{j}f),v^{*})\lambda_{a}^{n-j-1}V(x)$
as $n\to\infty$ for every fixed $j\geq 0$ and $x\in[0,1/(1-a)]$. Moreover, by (42), for every $n\geq 1$,
$\bigg|\lambda_{a}^{-n}P^{n}f(x)-\sum_{j=0}^{n-1}\frac{cV(x)}{\lambda_{a}^{j+1}}(M(P_{1}P_{2}^{j}f),v^{*})\bigg|\leq\lambda_{a}^{-n}\|P_{2}^{n}f\|+\sum_{j=0}^{n-1}\frac{C_{1}\gamma^{n}}{(\gamma\lambda_{a})^{j+1}}\|P_{1}P_{2}^{j}f\|.$ (70)
@@
$P_{2}f_{z}(x)=p\big(\mathbbm{1}\{ax+1>z\}-\mathbbm{1}\{1>z\}\big)+q{\mathbbm{1}}\{x\geq 1/a\}\big(\mathbbm{1}\{ax-1>z\}-\mathbbm{1}\{0>z\}\big).$
@@
$P_{1}P_{2}^{j}f_{z}(x)=pP_{2}^{j}f_{z}(1)\quad\text{for every }x\in[0,1/(1-a)],$ (72)
@@
$\sum_{j=0}^{\infty}\lambda_{a}^{-j-1}P_{1}P_{2}^{j}f_{z}(x)=\sum_{j=0}^{\varkappa_{a}(z)}\delta_{j}(z)(p/\lambda_{a})^{j+1}(q/p)^{L_{j}(z)}=:\overline{F}_{a}(z).$ (73)
Combining equalities (71) and (72) with estimate (70), where $(M\mathbbm{1},v^{*})=1$ for the constant function $\mathbbm{1}(x)\equiv 1$ by (38), and estimating the reminder of the sum in (73), we obtain
$\big|\lambda_{a}^{-n}P^{n}f_{z}(x)-c\overline{F}_{a}(z)V(x)\big|$ $\leq(C_{1}+\gamma\lambda_{a}/p)\gamma^{n}\sum_{j=0}^{\varkappa_{a}(z)\wedge n}(p/(\gamma\lambda_{a}))^{j+1}(q/p)^{L_{j}(z)}$
$\qquad+cV(x)\sum_{j=n}^{\varkappa_{a}(z)}(p/\lambda_{a})^{j+1}(q/p)^{L_{j}(z)}.$
Denote $r=\max(1,q/p)$ and $C_{2}^{\prime}=C_{1}+\gamma\lambda_{a}/p+cV(\tfrac{1}{1-a})$. Since $V(x)\leq V(\tfrac{1}{1-a})$, by Proposition 10 we get
$\big|\lambda_{a}^{-n}P^{n}f_{z}(x)-c\overline{F}_{a}(z)V(x)\big|$ $\leq C_{2}^{\prime}\gamma^{n}\sum_{j=0}^{n}(p/(\gamma\lambda_{a}))^{j+1}r^{C(j+1)+1}+C_{2}^{\prime}\sum_{j=n}^{\infty}(p/\lambda_{a})^{j+1}r^{C(j+1)+1}.$
This implies that for any $\gamma_{1}\in\big(\max(\gamma,pr^{C}/\lambda_{a}),1\big)$, there exists a constant $C_{2}>0$ such that
@@
for every $x,z\in[0,1/(1-a)]$ and $n\geq 1$. In particular, (74) implies that $\overline{F}_{a}(z)$ is finite.
@@
$\lim_{n\to\infty}\mathbb{P}_{x}(X_{n}>z|\tau>n)=\lim_{n\to\infty}\frac{P^{n}f_{z}(x)}{P^{n}\mathbbm{1}(x)}=\overline{F}_{a}(z)$ (75)
uniformly in $x,z\in[0,1/(1-a)]$ since $V(x)\geq V(0)\geq 1$. This limit does not depend on the starting point $x$. It also follows that the function $\overline{F}_{a}$ is non-increasing. Therefore, the conditional distributions converge weakly to the measure $\nu_{a}$ on $0,\frac{1}{1-a}]$ such that
@@
whenever $\overline{F}_{a}$ is continuous at $z$. This measure is a probability because the conditional distributions are tight, since they are supported on [0,1/(1a)][0,1/(1-a)].
6.2. Support and non-atomicity of $\nu_{a}$